Can you pls solve for a and b. I've tried and got a= -3, and b= 10, but the answer is a=10, b=3. I do not understand how they arrived at that answer.


The equations are: a - b = 7, and a - 4b = -7. Please solve for a and b and explain how you arrived at the answer. Thanks.
The Problem was taken from The Princeton Review, Math Smart.
ISBN 0-375-76216-7

Answer by checkley71(8403)

SORRY BUT A=10 & B=3 ARE NOT A SOLUTION.
a-b=7 or a=7+b
now replace a with (7+b) & solve for b.
7+b-4b=-7
-3b=-14
b=-14/-3
b=14/3 answer.
a-14/3=7
a=7+14/3
a=(21+14)/3
a=35/3 answer.
proof:
35/3-4(14/3)=-7
35/3-56/3=-7
-21/3=-7
-7=-7